OpenStudy (anonymous):

Help Please? Given the function f(x) = 2^x, find the value of f^−1(16).

OpenStudy (anonymous):

@hartnn I know you helped me with something like this before, but its different. Can you help again?

hartnn (hartnn):

sure

hartnn (hartnn):

what have you tried already ?

OpenStudy (anonymous):

I plugged 16 in as the x but im pretty sure that's not what you do...

hartnn (hartnn):

\(2^x =16 \)

OpenStudy (anonymous):

oh so I just do log2 on both sides?

hartnn (hartnn):

you can, but if you know \(16 = 2^{??}\) then it would be much simpler

hartnn (hartnn):

if you don't better take log on both sides

OpenStudy (anonymous):

what would be the simpler way?

hartnn (hartnn):

\(16 = 2^4 \) so ,in \(2^x = 16 \\ 2^x = 2^4 x= 4\)

hartnn (hartnn):

\(2^x = 16 \\ 2^x = 2^4 \\x= 4\)

OpenStudy (anonymous):

so f^-1(16)=4? and that is much simpler lol

hartnn (hartnn):

yes, it is :)

OpenStudy (anonymous):

okay thank you again :)

hartnn (hartnn):

welcome ^_^